# Ultra-stretchable and Self-Healable Vitrimers with Tuneable Damping and Mechanical Response

### v1

@@ Abstract
Vitrimers are a relatively new class of polymer materials with unique properties offered by cross-links that can undergo associative exchange dynamics. We here present a new class of vitrimers based on poly(methyl acrylate) with cross-links utilising dioxaboralane metathesis. These vitrimers demonstrate a combination of ultra-stretchability (up to 80 times their own length), mechanical toughness ( 40 MJ/m3), and thermal stability up to 250 °C; moreover, the vitrimers demonstrate excellent mechanical damping characterised by a loss factor () with a maximum of 2–3 and an effective value 0.3 across five decades in frequency (0.001–100 Hz), or correspondingly across a -range of 35 °C near room temperature (for a probe frequency of 1 Hz). The vitrimers can be successfully re-processed using both a thermo-mechanical and a chemical processing route, and can self-heal at room temperature, making them suitable for sustainable applications. The material properties are directly tuneable by variation of both the amount of cross-linker and by the degree of curing. Thus, this class of vitrimers are promising for applications where stretchability combined with mechanical toughness and/or a high mechanical dissipation is required.
@@ I Introduction
Elastomers that combine considerable stretchability with high mechanical strength and toughness are important for a wide range of applications including actuators, sensors, soft robotics, wearable electronics, and materials for energy storage. [1, 2, 3] Standard elastomers, such as vulcanised rubbers, can readily achieve tensile strains of several 100% combined with tensile strengths of 10 MPa. [4, 5, 6] However, elastomers that can sustain strains of several 1000% combined with a high tensile strength (a high mechanical toughness) are challenging to produce. A variety of approaches have been attempted often focused on combining an elastic stress-bearing network with a network of sacrificial bonds that can dissipate the applied mechanical energy by bond fracture. Reported polymer systems include: block-copolymer-based thermoplastic elastomers, [7] interpenetrating double polymer networks, [8, 9] elastomers based on supramolecular interactions, [10, 11] or polymeric networks based on dynamic covalent crosslinks, so called covalent adaptable networks (CANs) [12, 13, 14, 15, 16]. However, elastomers that show very high stretchability ( 1000–10000%), typically also exhibit significant strain-softening upon yielding, leading to a rapidly decreasing mechanical strength upon deformation. [10, 12, 14, 17, 18, 19, 20, 21] Thus, it remains an on-going challenge to produce elastomers that combine high stretchability, mechanical strength and toughness.
@@
Here, we present a new ultra-stretchable vitrimer system based on poly(methyl acrylate) (PMA), where the PMA-based vitrimers are prepared by free radical polymerization (FRP) of the methyl acrylate monomer and a diallyl dioxaborolane dynamic crosslinker. The associative crosslink exhange reaction, based on dioxaboralane metathesis, shows excellent thermal stability and a relatively low activation energy (16 kJ/mol), which allows rapid bond exchange without added solvent, catalyst or free diols. [31] Long chain-length PMAs are viscous fluids at room temperature, characterised by glass-transition temperatures 15°C, and are often used in coatings and adhesives, as liquid crystal elastomer backbone components, or as components in block co-polymers to improve material flexibility. [39, 40, 41] Here, PMA was chosen as the matrix polymer both due to its lack of crystallinity and due to its flexible nature, as reflected in a relatively low glass transition temperature, which ensures a high segmental mobility.
@@ II.1 Vitrimer synthesis and characterization
A set of poly(methyl acrylate) based vitrimers (VMAs) of varying crosslink densities were produced, as shown schematically in Figure 1(a); the dynamic crosslinker was first synthesized, [42] and subsequently copolymerized with the methyl acrylate monomer using FRP to produce the vitrimers. The schematic in Figure 1(b) illustrates how the associative bond exchange results in topological rearrangements, which in turn lead to malleability, even though the crosslink density is maintained. After synthesis, the samples were dried in an oven and then hot-pressed at °C, to produce vitrimers of nominal crosslink density and mol% (by stochiometry). 1H NMR spectra of the reactants (monomer and crosslinker) and the synthesis product are shown in Figure 1(c). The successful polymerization of methyl acrylate is demonstrated by the disappearance of signals due to the monomer double bond (peaks a-c in Figure 1(c)). The NMR spectra demonstrate a 50-60% incorporation of the crosslinker after the initial FRP step (peaks labelled 1-3 in Figure 1(c)). The lower (than 100%) incorporation is due to the lower reactivity of the double bond in the allyl compared to the acrylate moiety. [43] We refer to these first-cured samples as semi-cured and denote them as VMA-’ (the apostrophe marks that they are semi-cured and denotes the nominal molar cross-link density). To produce fully cured vitrimers, samples were further heated to °C for 2 hours under vacuum. After this process, no peaks due to the allyl double bond are observed in NMR, confirming that all vinyl bonds were fully incorporated; we thus refer to these samples as fully cured and denote them as VMA-. 1H NMR studies on the fully cured vitrimers (shown in Figure 2(a)) demonstrate that the resonance at 7.72 ppm, assigned to the phenyl ring of p-phenylene diboronic acid after hydrolysis (see inset), increases systematically with the nominal crosslink density, as expected for full incorporation of the crosslinker.
@@ II.2 Vitrimer Reprocessability
A strong advantage of vitrimers is the thermal reprocessability afforded by the dynamic nature of the cross-links [22, 31, 49]. Alternative chemical reprocessing routes, either by depolymerization or by degradation (e.g. through high temperature or acidic conditions), have also been demonstrated for vitrimers; [27, 50] however, these routes breaks down the vitrimer structure and reprocessing thus involves partially repeating the synthesis. In contrast, our vitrimers can be effectively re-processed using either a thermo-mechanical or a chemical route. For thermal reprocessing, the vitrimer samples were cut into pieces and placed into a hot-press mould, as shown in Figure 3(a)), where they were kept for 15 minutes at °C to facilitate dynamic cross-link exchange, and thus reprocessability. In contrast, the solvent-based reprocessing route is based on boronic ester hydrolysis, where the solvent needs to contain a sufficient amount of water for the hydrolysis (we used acetone mixed with a small amount of water; vol/vol = 200:1). The presence of the solvent initially leads to swelling of the vitrimers, followed by dissolution, where water acts to chemically “unzip” the cross-links (see the sketch in Figure 3(a)); NMR characterisation demonstrated that the vitrimer samples could be successfully hydrolysed using this approach, as shown in Figure 2(a). Following the dissolution process, the vitrimers were recovered by solvent removal in a vacuum oven at 80 °C, after which the dry samples were hot-pressed into samples of suitable dimensions. To investigate how successful each reprocessing route is, tensile stress-strain measurements were performed both before and after reprocessing. As shown in Figure 3(b) for VMA-1’, both the tensile strengths and elongation at break are largely recovered for samples undergoing either of the two reprocessing routes (2 MPa and 4500 %), and similar behaviour is observed for the corresponding fully cured sample (VMA-1), as shown in Figure S3.
@@ II.3 Vitrimer Network Response and Relaxation
The plateau modulus can generally be attributed both to the covalent dynamic crosslinks and to topological crosslinks due to chain entanglements. In Figure S4, we compare the crosslink density , as determined from the plateau modulus (), with the nominal crosslink density determined from the chemical composition, showing that and demonstrating an approximately linear relation, even though a significant difference is observed between the magnitude of the nominal and actual crosslink density. The average molecular weight between crosslinks () can be estimated from , [52] where is the mass density 0.8 g/cm3 at 160 °C for VMA (see the detailed discussion in the Experimental Section). The estimated for VMA- (1, 2.5, 5 and 8) are: 80 kDa, 14 kDa, 6 kDa, and 4 kDa, respectively. For comparison, the entanglement molecular weight () for linear poly(methyl acrylate) (PMA) is 11 kDa, [44], so that is larger than at low crosslink density, and smaller than at high crosslink density. For context, it is also important to compare both and with the length of the highly polydisperse primary chains given in Table S1. For VMA-1 and VMA-2.5 we find ; for VMA-5 and VMA-8 we find . In all cases, due to the high polydispersity, we thus expect a significant number of chains that have only a small number of dynamic crosslinks, and/or which are shorter than the entanglement molecular weight.
@@
Both with regards to the significant difference between and , and the small influence of entanglements in determining the ‘rubbery’ plateau, we note that: (i) since the reactivity of the crosslinker is much lower than that of the acrylate monomer, we expect a degree of heterogeneity in the cross-link distribution, and the formation of “effective cross-links” (each containing several nominal cross-links) is not surprising; (ii) the calculation of the nominal cross-link density does not take account of the possibility of cross-link formation within the same chain; such ‘loops’ are not generally stress-bearing and will thus not contribute to the plateau modulus; (iii) it is highly likely, especially at low crosslink density, that many of the crosslinks lie within branched structures that are either “free” chains wholly unconnected to the network, or “side branches” connected to the network at one point only: such structures can relax their internal configuration, and so relax their contribution to the stress, without breaking of crosslinks and thus do not contribute to the plateau; (iv) finally, relaxation of such free chains or side branches will release entanglements through the well known process of dynamic dilution [51], thus progressively increasing the effective entanglement molecular weight and reducing the influence of entanglements on the plateau.
@@
The relaxation times are shown in an Arrhenius plot in Figure 4(e). For each of the four samples, demonstrate Arrhenius behaviour: ( denotes a particular sample), where is the gas constant, as shown by the fits (dashed lines). The activation enthalpy (), determined for each of the four fits, are plotted in Figure 4(f) (squares). We find that is of similar magnitude 80 kJ/mol for the three higher crosslink density samples, whereas a somewhat lower of kJ/mol is found for the 1 mol % sample. These values are slightly higher than those previously reported for vitrimers based on the same dynamic crosslinker, where examples include: polysiloxane vitrimers (20–57 kJ/mol), [53] vitrimers prepared by a thiol-ene “click” reaction (50 kJ/mol), [54] polyolefin vitrimers (8–47 kJ/mol), [55] and poly(styrene-b-isoprene-b-styrene) vitrimers (44 kJ/mol). [56] As shown in Figure 4(f) the activation energy increases monotonically with increasing nominal crosslink density, even though the increase for crosslink densities above 2.5 mol % is very small; this behaviour is consistent with reports for vitrimers based on both dioxaboralane and other exchange mechanisms. [29, 57]
@@ II.4 Tensile deformation and stretchability
For the lowest applied strain rate, only a very weak strain hardening effect is found and a nearly horizontal plateau is observed, whereas higher rates lead to an increasing growth of the stress response. For an elastomer that is capable of large tensile strains, the engineering strain, which assumes a constant cross-sectional area, is not always representative of the tensile strength. Thus, in the inset to Figure 5(a) we also present the estimated true stress () by assuming a constant volume during stretching. For the three lowest extension speeds, the true stress at break is effectively constant at 80 MPa, whereas at the slowest deformation, the true stress reaches only 50 MPa – hence, deformation at a faster rate results in the vitrimer reaching its ‘true’ strength at break, and thus fracture, sooner.
@@
Macroscopic failure of materials commonly starts with micro-cracks, where crack propagation leads to the final failure. [69] For vitrimers, both a sufficient segmental mobility (low enough ) and any associative bond kinetics fast enough to allow network relaxations on the relevant deformation time-scale [61] could deter these mechanisms and thus contribute to large stretchability. Since °C is close to room temperature, the characteristic time-scale of the structural () relaxation is relatively slow, and 0.1 s (see Fig. S8). As an example for the four semi-cured samples, the imposed extensional strain rates () are 0.0267, 1.691, 2.251, and 3.024 s-1, respectively. In turn, the time-scales corresponding to the inverse applied strain rates () are 37.5, 0.59, 0.49, and 0.33 s, demonstrating that the time-scale characteristic of the deformation is for the three higher strain rates, whereas for the lowest applied rate. Thus, particularly for the slowest deformation rate, there is enough time for structural relaxation to take place during the deformation. It is also possible that the rate of the relaxation and/or the bond exchange mechanism is significantly sped up, particularly where the stress is focused near the tip of any micro-cracks, resulting in further contributions to these effects.
The overall result is a negligible strain-hardening for the lowest applied strain rate and a remarkably stretchable material, for which a deformation at break of 8000% is observed; the corresponding deformation at break is reduced to 2500% for the faster rate of 3.024 s-1, whereas the fracture strength is increased from 0.6 MPa to 3 MPa. 111We note that the rate of cross-link exchange in the quiescent vitrimers is relatively slow, as shown by the time-scale of the slow relaxation, as shown in Fig. S8; thus we expect that only if the rate of the cross-link exchange is driven by the applied deformation would it contribute to the stretchability.

### v2

@@ Abstract
Vitrimers are a relatively new class of polymer materials with unique properties offered by cross-links that can undergo associative exchange dynamics. We here present a new class of vitrimers based on poly(methyl acrylate) with cross-links utilising dioxaborolane metathesis. These vitrimers demonstrate a combination of ultra-stretchability (up to 80 times their own length), mechanical toughness ( 40 MJ/m3), and thermal stability up to 250 °C; moreover, the vitrimers demonstrate excellent mechanical damping characterised by a loss factor () with a maximum of 2–3 and an effective value 0.3 across five decades in frequency (0.001–100 Hz), or correspondingly across a -range of 35 °C near room temperature (for a probe frequency of 1 Hz). The vitrimers can be successfully re-processed using both a thermo-mechanical and a chemical processing route, and can for low crosslink density self-heal at room temperature, making them suitable for sustainable applications. The material properties are directly tuneable by variation of both the amount of cross-linker and by the degree of curing. Thus, this class of vitrimers are promising for applications where stretchability combined with mechanical toughness and/or a high mechanical dissipation is required.
@@ I Introduction
Elastomers that combine considerable stretchability with high mechanical strength and toughness are important for a wide range of applications including actuators, sensors, soft robotics, wearable electronics, and materials for energy storage. [1, 2, 3] Standard elastomers, such as vulcanised rubbers, can readily achieve tensile strains of several 100% combined with tensile strengths of 10 MPa. [4, 5, 6] However, elastomers that can sustain strains of several 1000% combined with a high tensile strength (a high mechanical toughness) are challenging to produce. A variety of approaches have been attempted often focused on combining an elastic stress-bearing network with a network of sacrificial bonds that can dissipate the applied mechanical energy by bond fracture. Reported polymer systems include: block-copolymer-based thermoplastic elastomers, [7] interpenetrating double polymer networks, [8, 9] elastomers based on supramolecular interactions, [10, 11, 12] or polymeric networks based on dynamic covalent crosslinks, so called covalent adaptable networks (CANs) [13, 14, 15, 16, 17]. However, elastomers that show very high stretchability ( 1000–10000%), typically also exhibit significant strain-softening upon yielding, leading to a rapidly decreasing mechanical strength upon deformation. [10, 13, 15, 18, 19, 20, 21, 22] Thus, it remains an on-going challenge to produce elastomers that combine high stretchability, mechanical strength and toughness.
@@
Here, we present a new ultra-stretchable vitrimer system based on poly(methyl acrylate) (PMA), where the PMA-based vitrimers are prepared by free radical polymerization (FRP) of the methyl acrylate monomer and a diallyl dioxaborolane dynamic crosslinker. The associative crosslink exchange reaction, based on dioxaborolane metathesis, shows excellent thermal stability and a relatively low activation energy (16 kJ/mol), which allows rapid bond exchange without added solvent, catalyst or free diols. [32] Long chain-length PMAs are viscous fluids at room temperature, characterised by glass-transition temperatures 15 °C, and are often used in coatings and adhesives, as liquid crystal elastomer backbone components, or as components in block co-polymers to improve material flexibility. [40, 41, 42] Here, PMA was chosen as the matrix polymer both due to its lack of crystallinity and due to its flexible nature, as reflected in a relatively low glass transition temperature, which ensures a high segmental mobility.
@@ II.1 Vitrimer synthesis and characterization
A set of poly(methyl acrylate) based vitrimers (VMAs) of varying crosslink densities were produced, as shown schematically in Figure 1(a); the dynamic crosslinker was first synthesized, [43] and subsequently copolymerized with the methyl acrylate monomer using FRP to produce the vitrimers. The schematic in Figure 1(b) illustrates how the associative bond exchange results in topological rearrangements, which in turn lead to malleability, even though the crosslink density is maintained. After synthesis, the samples were dried in an oven and then hot-pressed at °C, to produce vitrimers of nominal crosslink density and mol% (by stochiometry). 1H NMR spectra of the reactants (monomer and crosslinker) and the synthesis product are shown in Figure 1(c). The successful polymerization of methyl acrylate is demonstrated by the disappearance of signals due to the monomer double bond (peaks a-c in Figure 1(c)). The NMR spectra demonstrate a 50%–60% incorporation of the crosslinker after the initial FRP step (peaks labelled 1-3 in Figure 1(c)). The lower (than 100%) incorporation is due to the lower reactivity of the double bond in the allyl compared to the acrylate moiety. [44] We refer to these first-cured samples as semi-cured and denote them as VMA-’ (the apostrophe marks that they are semi-cured and denotes the nominal molar cross-link density). To produce fully cured vitrimers, samples were further heated to °C for 2 hours under vacuum. After this process, no peaks due to the allyl double bond are observed in NMR, confirming that all vinyl bonds were fully incorporated; we thus refer to these samples as fully cured and denote them as VMA-. 1H NMR studies on the fully cured vitrimers (shown in Figure 2(a)) demonstrate that the resonance at 7.72 ppm, assigned to the phenyl ring of p-phenylene diboronic acid after hydrolysis (see inset), increases systematically with the nominal crosslink density, as expected for full incorporation of the crosslinker.
@@ II.2 Vitrimer Reprocessability
A strong advantage of vitrimers is the thermal reprocessability afforded by the dynamic nature of the cross-links [23, 32, 50]. Alternative chemical reprocessing routes, either by depolymerization or by degradation (e.g. through high temperature or acidic conditions), have also been demonstrated for vitrimers; [28, 51] however, these routes breaks down the vitrimer structure and reprocessing thus involves partially repeating the synthesis. In contrast, our vitrimers can be effectively re-processed using either a thermo-mechanical or a chemical route. For thermal reprocessing, the vitrimer samples were cut into pieces and placed into a hot-press mould, as shown in Figure 3(a), where they were kept for 15 minutes at °C to facilitate dynamic cross-link exchange, and thus reprocessability. In contrast, the solvent-based reprocessing route is based on boronic ester hydrolysis, where the solvent needs to contain a sufficient amount of water for the hydrolysis (we used acetone mixed with a small amount of water; vol/vol = 200:1). The presence of the solvent initially leads to swelling of the vitrimers, followed by dissolution, where water acts to chemically “unzip” the cross-links (see the sketch in Figure 3(a)); NMR characterisation demonstrated that the vitrimer samples could be successfully hydrolysed using this approach, as shown in Figure 2(a). Following the dissolution process, the vitrimers were recovered by solvent removal in a vacuum oven at 80 °C, after which the dry samples were hot-pressed into samples of suitable dimensions. To investigate how successful each reprocessing route is, tensile stress-strain measurements were performed both before and after reprocessing. As shown in Figure 3(b) for VMA-1’, both the tensile strengths and elongation at break are largely recovered for samples undergoing either of the two reprocessing routes (2 MPa and 4500%), and similar behaviour is observed for the corresponding fully cured sample (VMA-1), as shown in Figure S3.
@@ II.3 Vitrimer Network Response and Relaxation
The plateau modulus can generally be attributed both to the covalent dynamic crosslinks and to topological crosslinks due to chain entanglements. The average effective molecular weight between crosslinks () can be estimated from , [53] where is the mass density, which we estimate using the literature value g/cm3 for linear poly(methyl acrylate) at the relevant reference temperature of 160 °C [54]. The resulting for VMA- (1, 2.5, 5 and 8) are: 109, 19, 8, and 5 kg/mol, respectively. For comparison, the entanglement molecular weight () for linear poly(methyl acrylate) (PMA) is 11 kg/mol, [45], so that is larger than at low crosslink density, and smaller than at high crosslink density. For context, it is also important to compare both and with the length of the highly polydisperse primary chains given in Table S1. For VMA-1 and VMA-2.5 we find ; for VMA-5 and VMA-8 we find . In all cases, due to the high polydispersity, we thus expect a significant number of chains that have only a small number of dynamic crosslinks, and/or which are shorter than the entanglement molecular weight.
Moreover, in Figure S4 we compare the estimated effective crosslink densities , as determined from the plateau moduli, with the corresponding nominal crosslink densities determined from the chemical composition and the mass density for the fully cured vitrimers. We determine using both the classical affine and phantom network models [53], respectively. Thus, the molar volumetric number density of stress-bearing chain strands was determined using , where for the affine network model, for the phantom network model, and is the cross-link functionality. We here use as an approximation that is consistent with the dynamic cross-linker chemistry [55], and for either model the cross-link density can be determined as . We find that for the affine network model the ratio lies within the range 0.04–0.12, and for the phantom network model within the range 0.08–0.25, for the investigated vitrimer samples. Thus, a significant difference between the estimated and the nominal crosslink density is consistently observed.
@@
Both with regards to the difference between and , and the small influence of entanglements in determining the ‘rubbery’ plateau, we note that: (i) since the reactivity of the crosslinker is much lower than that of the acrylate monomer, we expect a degree of heterogeneity in the cross-link distribution, and the formation of “effective cross-links” (each containing several nominal cross-links) is not surprising; (ii) the calculation of the nominal cross-link density does not take account of the possibility of cross-link formation within the same chain; such ‘loops’ are not generally stress-bearing and will thus not contribute to the plateau modulus; (iii) it is highly likely, especially at low crosslink density, that many of the crosslinks lie within branched structures that are either “free” chains wholly unconnected to the network, or “side branches” connected to the network at one point only: such structures can relax their internal configuration, and so relax their contribution to the stress, without breaking of crosslinks and thus do not contribute to the plateau; (iv) finally, relaxation of such free chains or side branches will release entanglements through the well known process of dynamic dilution [52], thus progressively increasing the effective entanglement molecular weight and reducing the influence of entanglements on the plateau.
@@
The relaxation times are shown in an Arrhenius plot in Figure 4(e). For each of the four samples, demonstrate Arrhenius behaviour: ( denotes a particular sample), where is the gas constant, as shown by the fits (dashed lines). The activation enthalpy (), determined for each of the four fits, are plotted in Figure 4(f) (squares). We find that is of similar magnitude 80 kJ/mol for the three higher crosslink density samples, whereas a somewhat lower of kJ/mol is found for the 1 mol% sample. These values are slightly higher than those previously reported for vitrimers based on the same dynamic crosslinker, where examples include: polysiloxane vitrimers (20–57 kJ/mol), [56] vitrimers prepared by a thiol-ene “click” reaction (50 kJ/mol), [57] polyolefin vitrimers (8–47 kJ/mol), [58] and poly(styrene-b-isoprene-b-styrene) vitrimers (44 kJ/mol). [59] As shown in Figure 4(f) the activation energy increases monotonically with increasing nominal crosslink density, even though the increase for crosslink densities above 2.5 mol% is very small; this behaviour is consistent with reports for vitrimers based on both dioxaborolane and other exchange mechanisms. [30, 55]
@@ II.4 Tensile deformation and stretchability
For the lowest applied strain rate, only a very weak strain hardening effect is found and a nearly horizontal plateau is observed, whereas higher rates lead to an increasing growth of the stress response. For an elastomer that is capable of large tensile strains, the engineering strain, which assumes a constant cross-sectional area, is not always representative of the tensile strength. Thus, in the inset to Figure 5(a) we also present the estimated true stress () by assuming a constant volume during stretching. For the three higher extension speeds, the true stress at break is effectively constant at 80 MPa, whereas at the slowest deformation, the true stress reaches only 50 MPa – hence, deformation at a faster rate results in the vitrimer reaching its ‘true’ strength at break, and thus fracture, sooner.
@@
Macroscopic failure of materials commonly starts with micro-cracks, where crack propagation leads to the final failure. [71] For vitrimers, both a sufficient segmental mobility (low enough ) and any associative bond kinetics fast enough to allow network relaxations on the relevant deformation time-scale [63] could deter these mechanisms and thus contribute to large stretchability. Since °C is close to room temperature, the characteristic time-scale of the structural () relaxation is relatively slow, and 0.1 s (see Fig. S8). As an example for the four semi-cured samples, the imposed extensional strain rates () are 0.0267, 1.691, 2.251, and 3.024 s-1, respectively. In turn, the time-scales corresponding to the inverse applied strain rates () are 37.5, 0.59, 0.49, and 0.33 s, demonstrating that the time-scale characteristic of the deformation is for the three higher strain rates, whereas for the lowest applied rate. Thus, particularly for the slowest deformation rate, there is enough time for structural relaxation to take place during the deformation. The overall result is a negligible strain-hardening for the lowest applied strain rate and a remarkably stretchable material, for which a deformation at break of 8000% is observed; the corresponding deformation at break is reduced to 2500% for the faster rate of 3.024 s-1, whereas the fracture strength is increased from 0.6 MPa to 3 MPa.
We note that it is possible that the rate of the relaxation and/or the bond exchange mechanism becomes significantly sped up during the deformation, particularly where the stress is focused near the tip of any micro-cracks, resulting in further contributions to the observed behaviour. The rate of cross-link exchange in the quiescent vitrimers is relatively slow, as shown by the time-scale of the network relaxation (see Fig. S8). Thus, we expect that only if this rate is significantly increased due to the applied deformation, will the dynamic nature of the crosslinks influence the stretchability. From our experiments alone, we cannot determine the role (if any) of the dynamic cross-links in the tensile deformation behaviour. However, previous studies on PMA-based elastomers with permanent crosslinks [72, 73] have also demonstrated highly stretchable behaviour, with a maximum stretchability of 1200% [72] observed for the lowest investigated crosslink density ( mol/cm3) and strain rate (0.01 s-1). This does not exclude a role of the dynamic cross-links, but it does illustrate the key role played by the near room temperature location of the glass transition for PMA in providing a highly stretchable material.